# From Uniform to Adaptive: General Skip-Block Mechanisms for Efficient PDE Neural Operators

### v1

@@ 4.1 Problem Formulation
In practice, we work with discretized representations of these functions. An input function $a$ is represented by its values at $N$ points in the physical domain, $\{a({\bm{x}}_{i})\}_{i=1}^{N}$. These $N$ points are treated as a set of tokens, which are then projected into a $d$-dimensional feature space to form an input matrix ${\bm{X}}_{\text{in}}\in\mathbb{R}^{N\times d}$. A standard, $L$-layer Transformer-based neural operator, denoted as $F(\cdot;{\bm{\theta}})$, can be formulated as a function that maps this input matrix to an output matrix ${\bm{Y}}_{\text{out}}\in\mathbb{R}^{N\times d}$ representing the solution field:
@@
where $f_{l}$ represents the $l$-th Transformer block parametrized by a subset of $\bm{\theta}$}, and $\circ$ denotes function composition. In this standard formulation, each block $f_{l}$ processes the full set of $N$ tokens, leading to the uniform computation bottleneck.
The central challenge we address is to reformulate this process. Instead of applying each $f_{l}$ to all $N$ tokens, our adaptive framework learns a policy to select a subset of $k_{l}$ active tokens ($k_{l}\leq N$) for each layer $l$. The computation at layer $l$ is then performed only on this active subset. The goal is to significantly reduce the total computational cost, which scales with $\sum_{l=1}^{L}k_{l}$, while preserving the predictive accuracy of the operator.
@@ 4.2 The Global Router Module
Formally, given the initial token embedding matrix ${\bm{X}}_{0}\in\mathbb{R}^{N\times d}$, where $N$ is the number of tokens and $d$ is the feature dimension, the router module $R$ processes this input to generate a vector of scalar importance scores. This importance score vector ${\bm{s}}\in\mathbb{R}^{N}$ is computed as follows:
@@ C.2 Quantile-Based Region Partitioning
After computing the complexity scores $\{s_{1},s_{2},\dots,s_{N}\}$ for all $N$ points in a sample, we partition the domain into regions via quantile-based thresholds. This ensures that the division adapts dynamically to different samples.

### v2

@@ 4.1 Problem Formulation
In practice, we work with discretized representations of these functions. An input function $a$ is represented by its values at NN points in the physical domain, $\{a({\bm{x}}_{i})\}_{i=1}^{N}$. These NN points are treated as a set of tokens, which are then projected into a $d$-dimensional feature space to form an input matrix ${\bm{X}}_{\text{in}}\in\mathbb{R}^{N\times d}$. A standard, $L$-layer Transformer-based neural operator, denoted as $F(\cdot;{\bm{\theta}})$, can be formulated as a function that maps this input matrix to an output matrix ${\bm{Y}}_{\text{out}}\in\mathbb{R}^{N\times d}$ representing the solution field:
@@
where $f_{l}$ represents the $l$-th Transformer block parametrized by a subset of $\bm{\theta}$}, and $\circ$ denotes function composition. In this standard formulation, each block $f_{l}$ processes the full set of NN tokens, leading to the uniform computation bottleneck.
The central challenge we address is to reformulate this process. Instead of applying each $f_{l}$ to all NN tokens, our adaptive framework learns a policy to select a subset of $k_{l}$ active tokens ($k_{l}\leq N$) for each layer $l$. The computation at layer $l$ is then performed only on this active subset. The goal is to significantly reduce the total computational cost, which scales with $\sum_{l=1}^{L}k_{l}$, while preserving the predictive accuracy of the operator.
@@ 4.2 The Global Router Module
Formally, given the initial token embedding matrix ${\bm{X}}_{0}\in\mathbb{R}^{N\times d}$, where NN is the number of tokens and $d$ is the feature dimension, the router module $R$ processes this input to generate a vector of scalar importance scores. This importance score vector ${\bm{s}}\in\mathbb{R}^{N}$ is computed as follows:
@@ C.2 Quantile-Based Region Partitioning
After computing the complexity scores $\{s_{1},s_{2},\dots,s_{N}\}$ for all NN points in a sample, we partition the domain into regions via quantile-based thresholds. This ensures that the division adapts dynamically to different samples.